What is the pH of a saturated solution of Al(OH)3? For Al(OH)3 , Ksp=2.0x10-33.
Answer #1

Ksp for Al(OH)2 = (Al+3][OH-]3 = 2.0 x 10-33 Let S be the solubility of Al(OH)3 . Al(OH)3 = Alt + 3OH- 35 solid S Ksp = (S)(3